HRMS, 28th Colloquium, Dijon, France, Aug 28 - Sept 1, 2023

PROGRESS IN THE ZURICH EXPERIMENT ON PARITY VIOLATION IN CHIRAL MOLECULES

\underline{GUNTHER WICHMANN} , KAREN KEPPLER, GEORG SEYFANG,

MARTIN QUACK,

Department of Chemistry and Applied Biosciences, ETH Zurich, 8093 Zurich, Switzerland, www.ir.ETHz.ch

Because of the space inversion symmetry of the electromagnetic force, the ground state energies of the enantiomers of chiral molecules would be identical by symmetry in conventional quantum chemical theory. Including also the parity violating weak nuclear force, they are, however, predicted to be different by a very small parity violating energy difference Δ_{pv}E in the sub-feV and low feV range. This small effect has so far never been measured and such experiments constitute a major challenge of physical-chemical stereochemistry with possible importance both for the fundamental physics in the standard model of elementary particles (SMPP) and for the evolution of biomolecular homochirality in the origin of life as reviewed in [1, 2]. An important step in our approach following [3] has been the proof of concept testing the experimental stability of parity states with the achiral molecule NH_{3} [4] demonstrating the possibility of detecting values of Δ_{pv}E as small as 100 aeV (or larger). An important next step is to find a parity state in a chiral molecule by measuring and analyzing mid-IR vibration-rotation-tunneling spectra (around 3000 cm^{−1}) of one of our mid-sized chiral candidate molecules with low vapor pressure (listed e.g. in Tab. 2of Ref. [2]). Here, we report improved cw-laser cavity ring-down spectroscopy [5, 6, 7] referenced to a frequency comb combined with a slit jet expansion and partial pressures below 1 mbar. We report the spectrum of benzene spanning ranges of more than 10 GHz and other molecules when available at the time of this conference in the mid-IR with complexity similar to the spectroscopy of our chiral candidate molecule, 1,2-dithiine C_{4}H_{4}S_{2} [8, 9], see also the work on 1,3-difluoroallene CHFCCHF [10, 11] and ClSSCl [12]. References

[1] M. Quack, G. Seyfang, G. Wichmann, Proc. Nobel Symposium 167 on ’Chiral Matter’ (Stockholm 2021),

World Scientific Singapore, 2023, 209-268.

[2] M. Quack, G. Seyfang, G. Wichmann, Chem. Sci., 13, 2022, 10598-10643.
[3] M. Quack, Chem. Phys. Letters., 132, 1986, 147-153.
[4] P. Dietiker, E. Milogyadov, M. Quack, A. Schneider, and G. Seyfang, J. Chem. Phys., 143, 2015, 244305.
[5] M. Snels, V. Horká-Zelenková, H. Hollenstein and M. Quack, in Handbook of High−resolution Spectroscopy

edited by M. Quack and F. Merkt (Wiley, Chichester, New York, 2011), vol. 2, pp. 1021-1067.

[6] Y. B. He, M. Hippler and M. Quack, Chem. Phys. Lett., 289, 1998, 527-534.
[7] G. Wichmann, E. Miloglyadov, G. Seyfang and M. Quack, Mol. Phys., 118, 2020, e1752946.
[8] S. Albert, I. Bolotova, Z. Chen, C. Fábri, L. Horný, M. Quack, G. Seyfang and D. Zindel,

Phys. Chem. Chem. Phys., 18, 2016, 21976-21993.

[9] S. Albert, F. Arn, I. Bolotova, Z. Chen, C. Fábri, G. Grassi, P. Lerch, M. Quack, G. Seyfang, A. Wokaun,

D. Zindel, J. Phys. Chem. Lett., 7, 2016, 3847-3853.

[10] M. Gottselig, M. Quack, J. Chem. Phys., 123, 2005, 84305-1 - 84305-11.

[11] Keppler et al. Poster at this meeting.

[12] G. Wichmann, G. Seyfang, and M. Quack, Mol. Phys., 119, 2021, e1959073
p-number: p110  Submitted on Tue, 13 Jun 2023 17:58:16 +0200